### v2

# One-Step Diffusion for Detail-Rich and Temporally Consistent Video Super-Resolution

Abstract
It is a challenging problem to reproduce rich spatial details while maintaining temporal consistency in real-world video super-resolution (Real-VSR), especially when we leverage pre-trained generative models such as stable diffusion (SD) for realistic details synthesis. Existing SD-based Real-VSR methods often compromise spatial details for temporal coherence, resulting in suboptimal visual quality. We argue that the key lies in how to effectively extract the degradation-robust temporal consistency priors from the low-quality (LQ) input video and enhance the video details while maintaining the extracted consistency priors. To achieve this, we propose a Dual LoRA Learning (DLoRAL) paradigm to train an effective SD-based one-step diffusion model, achieving realistic frame details and temporal consistency simultaneously. Specifically, we introduce a Cross-Frame Retrieval (CFR) module to aggregate complementary information across frames, and train a Consistency-LoRA (C-LoRA) to learn robust temporal representations from degraded inputs. After consistency learning, we fix the CFR and C-LoRA modules and train a Detail-LoRA (D-LoRA) to enhance spatial details while aligning with the temporal space defined by C-LoRA to keep temporal coherence. The two phases alternate iteratively for optimization, collaboratively delivering consistent and detail-rich outputs. During inference, the two LoRA branches are merged into the SD model, allowing efficient and high-quality video restoration in a single diffusion step. Experiments show that DLoRAL achieves strong performance in both accuracy and speed. Code and models are available at https://github.com/yjsunnn/DLoRAL.
[…]
1 Introduction
Video super-resolution (VSR) aims to reconstruct high-quality (HQ) videos from low-quality (LQ) inputs. Traditional VSR methods typically rely on convolutional neural network (CNN)-based wang2019edvr ; chan2022basicvsr++ and Transformer-based designs cao2021vrt ; liu2022learning , trained with pixel-wise $L_{2}$ or $L_{1}$ losses. While effective in some metrics (e.g., PSNR), these methods often produce over-smoothed results without fine details. To improve perceptual quality, generative adversarial network (GAN)-based VSR methods incorporate the adversarial loss ledig2017photo during training to encourage sharper details restoration bulat2018superfan ; menon2020pulse ; bulat2018learnagan ; xu2024videogigagan . However, many VSR models xue2019TOF ; fuoli2019duf ; zhang2024tmp are trained under simplified degradation assumptions (e.g., bicubic downsampling), limiting their performance on real-world LQ videos with complex and unknown degradations. Additionally, GAN-based methods can produce unnatural artifacts and generalize poorly to diverse video content. Recently, pre-trained diffusion-based text-to-image (T2I) models such as Stable Diffusion (SD) rombach2022stablediffusion ; blattmann2023stablevideodiffusion have shown impressive results in real-world image super-resolution (Real-ISR) wang2024stablesr ; wu2024seesr ; yu2024supir ; sun2023improving ; wu2024osediff ; sun2024pisa ; wang2024sinsr with realistic textures. One line of research treats the LQ image as a control signal and employs ControlNet-like structures zhang2023controlnet to guide generation wang2024stablesr ; yu2024supir ; wu2024seesr ; sun2023improving , and another line of research directly fine-tunes the SD model with LoRA hu2022lora for efficient one-step restoration wu2024osediff ; sun2024pisa .
The success of SD in Real-ISR inspired exploration of diffusion models for real-world video super-resolution (Real-VSR). Although the powerful generative priors of SD can enhance details, they can introduce inconsistencies among frames when the generated textures sometimes deviate from the content of the LQ inputs wang2024stablesr ; sun2024pisa . To alleviate this issue, existing SD-based Real-VSR methods typically suppress such fluctuations at the cost of perceptual quality. These methods, such as Upscale-A-Video zhou2024upscale and MGLD-VSR yang2024motion , incorporate temporal modules into pre-trained SD models and adopt frame-wise losses to balance spatial detail and temporal consistency. Despite the significant progress achieved, these methods have two major limitations. First, these approaches optimize detail and consistency jointly in a single model, resulting in suboptimal trade-offs. Improving one objective usually harms the other due to their conflicting nature. Second, the temporal consistency existing in real-world LQ videos is ignored, which can be effectively leveraged to help anchor detail generation on a consistent temporal basis.
(a) Quality comparison
(b) Efficiency Comparison
Figure 1: Quality and efficiency comparison among SD-based Real-VSR methods. (a) Quality comparison on the VideoLQ benchmark chan2022realbasicvsr . (b) Efficiency comparison tested on an A100 GPU ($512\times 512$ input with 50 frames for ×$\times 4$ VSR). DLoRAL achieves the best perceptual quality with only one diffusion step, about 10$\times$ faster than Upscale-A-Video zhou2024upscale , MGLD yang2024motion , and STAR xie2025star .
To address these issues, we propose a Dual LoRA Learning (DLoRAL) framework for Real-VSR. Our method is built on a one-step residual diffusion model wu2024osediff ; sun2024pisa , which significantly reduces inference time while maintaining strong generative capability. Inspired by PiSA-SR sun2024pisa , which learns two LoRA modules to achieve adjustable Real-ISR results, we design two decoupled LoRA branches within the shared diffusion UNet to resolve the conflict between spatial detail and temporal coherence. Specifically, a Consistency-LoRA (C-LoRA) is designed to learn temporal consistency representation, and a Detail-LoRA (D-LoRA) is designed to restore high-frequency spatial details. To exploit the inherent temporal consistency in LQ videos, we introduce a Cross-Frame Retrieval (CFR) module, which extracts structure-aligned temporal features from adjacent degraded frames, helping the model learn degradation-robust representations. CFR not only provides a stable and informative intermediate representation for C-LoRA to build upon, but also serves as the anchor for the subsequent detail enhancement stage to maintain temporal alignment.
Instead of optimizing both objectives jointly, we adopt a dual-stage training strategy. The training begins from the temporal consistency stage, in which we fine-tune C-LoRA and CFR modules using consistency-related losses. In the detail enhancement stage, we freeze C-LoRA and CFR, and train D-LoRA to refine high-frequency details with the additional classifier score distillation (CSD) sun2024pisa loss. These two stages are alternatively trained to allow each branch to specialize in its objective. During inference, the two LoRA modules can be integrated in one-step diffusion. As illustrated in Fig. 1, our DLoRAL method achieves both high temporal consistency and superior visual quality, outperforming previous Real-VSR methods in overall quality, as well as inference speed (about 10$\times$ speedup over current methods zhou2024upscale ; yang2024motion ; xie2025star , as illustrated in Fig. 1(b)).
Our main contributions are summarized as follows. (1) We propose a Dual LoRA Learning (DLoRAL) paradigm for Real-VSR, which decouples the learning of temporal consistency and spatial details into two dedicated LoRA modules under a unified one-step diffusion framework. (2) We introduce a Cross-Frame Retrieval (CFR) module to extract degradation-robust temporal priors for Consistency-LoRA (C-LoRA) training, providing structure-aligned intermediate representations that guide the subsequent training of Detail-LoRA (D-LoRA) for high-fidelity restoration. (3) Our DLoRAL model achieves state-of-the-art performance on Real-VSR benchmarks, producing visually realistic frame details and stable temporal consistency.
2 Related Work
Real-World VSR. Conventional VSR methods wang2019edvr ; fuoli2019duf ; li2020mucan typically rely on simply synthesized data (e.g., bicubic downsampling), leading to a significant performance gap when applied to real-world videos. Early works yang2021realvsr ; wei2023realbsr addressed this by collecting real-world LQ-HQ video pairs, such as the iPhone-captured dataset yang2021realvsr . However, these datasets are limited by device bias and scalability. The following works wang2021realesrgan ; chan2022realbasicvsr simulated realistic degradations by combining blur, noise, and compression, while others enhanced robustness through architectural design. For instance, RealVSR yang2021realvsr introduces a domain adaptation mechanism that aligns feature distributions between synthetic and real domains through adversarial learning. RealBasicVSR chan2022realbasicvsr proposes a degradation modeling framework that refines the restoration process through iterative correction modules. Despite these advances, existing methods still struggle to recover fine details and generalize across diverse real-world scenarios, often producing over-smoothed outputs.
Diffusion Based Real-VSR. Recent advances in diffusion models for image restoration ai2024dreamclear ; dong2024tsd ; qu2024xpsr ; yu2024scaling ; yue2024arbitrary have inspired the extension to Real-VSR tasks zhou2024upscale ; yang2024motion ; li2025diffvsr ; xie2025star . A common approach is to adapt pre-trained T2I models by injecting temporal modules to ensure both perceptual quality and temporal consistency. For example, Upscale-A-Video zhou2024upscale integrates temporal layers into the pre-trained diffusion model and proposes a flow-guided recurrent latent propagation module. MGLD-VSR yang2024motion guides the diffusion process with a motion-guided loss and inserts a temporal module into the diffusion decoder. The other directions include decomposing the complex learning burden into staged training phases li2025diffvsr and reformulating attention mechanisms in diffusion transformers wang2025seedvr to process videos of arbitrary length. Rather than leveraging the pre-trained T2I model, STAR xie2025star leverages compressed temporal representations from text-to-video (T2V) models blattmann2023stablevideodiffusion .Despite these efforts, balancing spatial detail and temporal consistency remains a key challenge. Most existing methods enforce frame-level constraints to improve consistency by sacrificing visual fidelity. In this work, we propose a decoupled learning strategy: first learning degradation-robust temporal priors from LQ inputs then guiding HQ generation with these features. This design ensures both high-quality detail restoration and stable temporal coherence.
Real-VSR Paradigms. Recent VSR methods follow two main paradigms: sliding-window-based wang2019edvr ; fuoli2019duf ; xue2019TOF ; li2020mucan and recurrent-based chan2021basicvsr ; chan2022basicvsr++ ; cao2021vrt ; liu2022TTVSR ; zhou2024upscale ; yang2024motion ; xie2025star . Sliding-window-based methods reconstruct each output frame using a set of neighboring frames, capturing fine-grained local details and short-term temporal dependencies. In contrast, recurrent-based methods propagate features across frames sequentially, offering higher efficiency, but are prone to error accumulation and detail degradation. Most diffusion-based Real-VSR methods zhou2024upscale ; yang2024motion ; xie2025star adopt the recurrent design for its inference efficiency. In this work, we build on a sliding-window framework to better preserve spatial and temporal details. To mitigate the computational overhead, we adopt a one-step diffusion strategy that eliminates redundancy while maintaining high reconstruction quality.
3 Methodology
3.1 Preliminary
Diffusion models rombach2022stablediffusion simulate a forward process where a clean latent code $z_{0}$ is gradually noised into $z_{t}$ using Gaussian noise: ${z_{t}}=\sqrt{{{\bar{\alpha}}_{t}}}\cdot{z_{0}}+\sqrt{1-{{\bar{\alpha}}_{t}}}\cdot\epsilon$, with $\epsilon\sim\mathcal{N}(0,I)$ and $\bar{\alpha}_{t}$ following a predefined schedule. During training, a model $\epsilon_{\theta}(t,z_{t})$ is trained to predict the added noise at each timestep $t$. During inference, $z_{0}$ is recovered from pure noise $z_{T}\sim\mathcal{N}(0,I)$ by iterative denoising. However, this multi-step process is slow and stochastic, limiting its efficiency and stability in super-resolution (SR) tasks that demand fast and reliable reconstruction. To address this issue, recent works  wu2024osediff ; sun2024pisa propose one-step diffusion that skips iterative sampling by directly refining an LQ latent into its HQ counterpart. To further improve controllability, PiSA-SR sun2024pisa introduces a residual learning formulation that allows the model to focuses on high-frequency corrections:
${z}^{HQ}=z^{LQ}-\epsilon_{\theta}(z^{LQ})$ (1)
where $z^{LQ}$ and $z^{HQ}$ represent the latent codes of LQ and HQ respectively.
Most existing VSR methods zhou2024upscale ; yang2024motion ; xie2025star rely on multi-step diffusion, resulting in high computational cost. In this work, we make the first attempt to apply a one-step diffusion framework to VSR, improving efficiency while preserving restoration quality by adapting the residual learning formulation in Eq. (1) to accelerate convergence. To this end, we introduce VSR-specific modules and learning strategies to produce detail-rich and temporally consistent results.
3.2 Dual LoRA Learning Network for Real-VSR
Motivation. There is a fundamental challenge in Real-VSR: how to balance the preservation of spatial details and the enforcement of temporal consistency. To simultaneously achieve both objectives, we begin by analyzing the characteristics of real-world LQ videos and the limitations of current SD-based VSR methods, which motivate the design of our proposed framework.
  • Temporal Consistency in Degraded Videos. Despite degradations, such as noise, blur, and compression, real-world LQ videos retain stable information across frames, preserving inherent structural and semantic consistency. Leveraging these consistent representations provides a strong foundation for reconstructing HQ videos with realistic details and temporal coherence. To exploit this, we propose a Cross-Frame Retrieval (CFR) module to aggregate complementary information across frames to enhance consistency. In addition, we design a Consistency-LoRA (C-LoRA) to further improve reconstruction by reinforcing temporal alignment and structural integrity. This stage lays the groundwork for more accurate guidance in the subsequent detail enhancement phase.
  • Conflict in Optimizing Details and Consistency. To adapt pre-trained diffusion models for VSR and balance spatial detail and temporal coherence, existing methods rombach2022stablediffusion ; yang2024cogvideox typically introduce trainable layers optimized jointly with diffusion and temporal losses. However, detail generation and consistency preservation are inherently conflicting objectives, and joint optimization often results in suboptimal trade-offs. To address this, we propose two decoupled weight spaces: one for temporal consistency modeling and another for detail enhancement. Rather than training two networks, which is costly, we adopt a decoupled scheme inspired by PiSA-SR sun2024pisa , embedding two specialized LoRA branches into a shared SD UNet. This lightweight design enables alternative refinement, allowing each branch to focus on its objective.
Framework Overview. Building on the above insights, we design a Dual LoRA Learning (DLoRAL) framework to generate HQ video outputs from degraded inputs. Given an LQ sequence of $N$ frames, $\mathbf{I}^{LQ}=\{I^{LQ}_{n}\mid n=1,\dots,N\}$, our model $G_{\theta}$ generates a corresponding HQ sequence $\mathbf{I}^{HQ}=\{I^{HQ}_{n}\mid n=1,\dots,N\}$. To utilize information from neighboring frames, we adopt a sliding-window strategy wang2019edvr ; fuoli2019duf ; li2020mucan , where each HQ frame $I^{HQ}_{n}$ is generated from two adjacent LQ frames: the current $n$-th frame $I^{LQ}_{n}$ and its preceding frame $I^{LQ}_{n-1}$. For the first frame $I_{1}^{LQ}$, which lacks a previous frame, we adopt a self-replication approach to generate $I_{1}^{LQ}$.
As illustrated in Fig. 2, our generator $G_{\theta}$ leverages the pre-trained SD model, which consists of a VAE encoder $E_{\theta}$, an SD UNet $\epsilon_{\theta}$, and a VAE decoder $D_{\theta}$. Our DLoRAL framework employs two specialized training stages, i.e., temporal consistency stage and detail enhancement stage. In the temporal consistency stage, the CFR module retrieves inter-frame relevant information from degraded inputs, then the UNet is finetuned by C-LoRA for further reinforcement of temporal alignment. In the detail enhancement stage, D-LoRA is optimized to improve spatial visual quality. These two stages are trained alternately in an iterative manner to progressively refine both temporal consistency and spatial quality, ultimately leading to coherent and detail-preserved video restoration. During inference, the C-LoRA and D-LoRA are merged into the SD UNet to ensure efficient deployment.
Figure 2: The training pipeline of our proposed DLoRAL. The Cross-Frame Retrieval (CFR) and Consistency-LoRA (C-LoRA) modules are optimized in the consistency stage, while the Detail-LoRA (D-LoRA) is optimized in the enhancement stage. Both stages are alternately trained to ensure temporal coherence and visual quality.
Temporal Consistency Stage. This stage is to establish a temporally coherent and robust representation from the LQ video sequence $\textbf{I}^{LQ}$ before enhancing details. This stage involves two main steps: temporal feature fusion using a CFR module and fine-tuning the SD UNet to improve consistency.
To unlock the inherent consistency among degraded inputs, CFR improves the current latent representation $z_{n}^{LQ}$ by employing a specialized attention mechanism that integrates complementary information from the previous latent feature $z_{n-1}^{LQ}$. Specifically, with encoded features $z_{n}^{LQ}$ and $z_{n-1}^{LQ}$, CFR first warps them into the same coordinate space with SpyNet ranjan2017spynet following a common frame alignment procedure wang2019edvr ; yang2024motion (denoted as $F_{wp}$). The current latent features $z_{n}^{LQ}$ and aligned latent features $F_{wp}(z_{n-1}^{LQ})$ are then projected into query ($Q_{n}$), key ($K_{n-1}$), and value ($V_{n-1}$) embeddings through $1\times 1$ convolutions (denoted as $\circ$) parameterized by $W_{Q}$, $W_{K}$, and $W_{V}$, as shown below:
$Q_{n}=W_{Q}\circ z^{LQ}_{n},\quad K_{n-1}=W_{K}\circ F_{wp}(z^{LQ}_{n-1}),\quad V_{n-1}=W_{V}\circ F_{wp}(z^{LQ}_{n-1}).$ (2)
With obtained embeddings extracted from adjacent frames, CFR employs two mechanisms to enhance fusion quality. First, for each query position $p$, it selectively attends to only the top-k most similar positions (denoted as $F_{topk}[p]$) in the aligned previous frame, avoiding perturbations from uncorrelated noises. Second, for each query position $p$, a learnable threshold $\tau_{n}[p]$ is predicted via a lightweight MLP. It dynamically adapts to regional characteristics - enforcing stricter filtering in detail-rich areas while being more permissive in flat regions, ensuring that only confident matches could contribute to the final fusion. The fused feature $\bar{z}^{LQ}_{n}[p]$ is computed as:
$\bar{z}^{LQ}_{n}[p]=z^{LQ}_{n}[p]+\sum_{q\in F_{topk}[p]}\phi\left(\frac{\langle Q_{n}[p],K_{n-1}[q]\rangle}{\sqrt{d}}-\tau_{n}[p]\right)\cdot V_{n-1}[q],$ (3)
where $\phi(\cdot)$ is a non-negative gating function (e.g., ReLU glorot2011ReLU ), and $d$ is the channel dimension.
The latent feature $\bar{z}^{LQ}_{n}$ is then processed by the UNet to generate the HQ latent $z^{HQ}_{n}$. In this stage, only the C-LoRA is trainable, while D-LoRA remains frozen. The final HQ frame is reconstructed via the VAE decoder by $I^{HQ}_{n}=D_{\theta}(z^{HQ}_{n})$. All trainable components in this stage, including the CFR module and C-LoRA, are optimized using the consistency loss $\mathcal{L}_{\text{cons}}$, which is designed to ensure both the quality of individual frames and the temporal consistency across the sequence. It combines the pixel-level loss ($\mathcal{L}_{\text{pix}}$), LPIPS loss ($\mathcal{L}_{\text{lpips}}$), and optical flow loss ($\mathcal{L}_{\text{opt}}$), as shown below:
$\begin{array}[]{l}\mathcal{L}_{\text{cons}}=\lambda_{\text{pix}}\mathcal{L}_{\text{pix}}+\lambda_{\text{lpips}}\mathcal{L}_{\text{lpips}}+\lambda_{\text{opt}}\mathcal{L}_{\text{opt}},\\[6.0pt] \mathcal{L}_{\text{opt}}=\left\|{O}^{HQ}_{n}-O_{n}^{\text{GT}}\right\|_{1}=\left\|F(I^{HQ}_{n},I^{HQ}_{n+1})-F(I_{n}^{\text{GT}},I_{n+1}^{\text{GT}})\right\|_{1}.\\ \end{array}$ (4)
Here, the $\ell_{2}$ loss is adopted as the $\mathcal{L}_{\text{pix}}$, and $\mathcal{L}_{\text{opt}}$ measures the $L_{1}$ distance between optical flow maps estimated from generated and ground-truth frame pairs, promoting motion alignment and smooth transitions. The loss weights $\lambda_{\text{pix}}$, $\lambda_{\text{lpips}}$, and $\lambda_{\text{opt}}$ are empirically set to balance spatial accuracy, perceptual quality, and temporal consistency.
Detail Enhancement Stage. Different from the temporal consistency stage, which yields aligned and coherent latent representations, the detail enhancement stage focuses on restoring high-frequency visual details. In this stage, adjacent latent features $z_{n-1}^{LQ}$ and $z_{n}^{LQ}$ are processed by the frozen CFR module to reapply the learned alignment and fusion, thus the temporal consistency learned in the consistency stage is maintained without introducing new variations.
The resulting temporally enriched latent $\bar{z}^{LQ}_{n}$ is then fed into the diffusion UNet $\epsilon_{\theta}$. We employ a decoupled finetuning strategy: only the D-LoRA parameters, responsible for detail synthesis, are trainable, while the C-LoRA parameters, associated with consistency, remain frozen. This setting allows the D-LoRA to focus solely on detail synthesis without compromising the temporal structure previously established. The output HQ latent $z^{HQ}_{n}$ is then decoded using the frozen decoder $D_{\theta}$ to produce the final super-resolved frame $I^{HQ}_{n}$.
To guide this detail enhancement while preserving the structure learned previously, the loss function $\mathcal{L}_{\text{enh}}$ combines several components as follows:
$\mathcal{L}_{\text{enh}}=\lambda_{\text{pix}}\mathcal{L}_{\text{pix}}+\lambda_{\text{lpips}}\mathcal{L}_{\text{lpips}}+\lambda_{\text{opt}}\mathcal{L}_{\text{opt}}+\lambda_{\text{csd}}\mathcal{L}_{\text{csd}}.$ (5)
We retain $\mathcal{L}_{\text{pix}}$, $\mathcal{L}_{\text{lpips}}$, and $\mathcal{L}_{\text{opt}}$ used in the consistency stage (as Eq. (4)), serving as anchors to maintain spatial fidelity and motion coherence. Furthermore, we introduce the Classifier Score Distillation (CSD) loss sun2024pisa , $\mathcal{L}_{\text{csd}}$, which encourages the generation of richer and finer details.
3.3 Training and Inference
Dynamic Dual-Stage Training. We adopt a dynamic dual-stage training scheme. The training begins with the consistency stage, aiming at learning degradation-robust features and establishing strong temporal coherence among frames. In this stage, only the CFR and C-LoRA modules are trainable, while the D-LoRA is fixed. Once the model converges in the consistency stage, the training switches to refine high-frequency spatial details, guided by $\mathcal{L}_{\text{enh}}$ with the additional CSD loss. In this stage, only the D-LoRA parameters are trainable, while the CFR module and C-LoRA are fixed. Such an alternative training is iterated, allowing the model to dynamically converge toward a solution that balances temporal coherence and visual fidelity.
Smooth Transition Between Training Stages. Compared to the consistency stage, the enhancement stage introduces an additional loss function $\mathcal{L}_{\text{csd}}$ for enriching semantic details. Directly switching between the full loss functions $\mathcal{L}_{\text{cons}}$ and $\mathcal{L}_{\text{enh}}$ can lead to instability due to the abrupt change in learning targets. To prevent this, we employ a re-weighting strategy that progressively shifts the loss objective, ensuring a smooth transition between stages. Taking the transition from the consistency stage to the enhancement stage as an example, after the consistency stage, the two loss functions are interpolated as the optimization objective for a warm-up phase of $s_{t}$ steps, as shown below:
$\mathcal{L}(s)=(1-\frac{s}{s_{t}})\cdot\mathcal{L}_{\text{cons}}+\frac{s}{s_{t}}\cdot\mathcal{L}_{\text{enh}},\quad s\in[0,s_{t}],$ (6)
where $s$ denotes the current step within the transition. Symmetric interpolation is applied when we switch back from the enhancement stage to the consistency stage.
Inference Phase. At test time, both C-LoRA and D-LoRA are activated and merged into the frozen diffusion UNet. A single diffusion step is used to enhance the LQ input to HQ video frames.
4 Experiment
4.1 Experimental Settings
Implementation Details. We adopt the pre-trained Stable Diffusion V2.1 as the backbone of denoising U-Net. Training is carried out with a batch size of 16, a sequence length of 3, and a video resolution of $512\times 512$. All models are trained using the PyTorch framework on 4 NVIDIA A100 GPUs. We use Adam optimizer with an initial learning rate of $5\times 10^{-5}$. For inference, both C-LoRA and D-LoRA are activated simultaneously in a frozen UNet. Videos are processed in sliding sequences to fit GPU memory limits.
Training Datasets. To support the decoupled training design of our DLoRAL framework, we construct two training datasets for the consistency and enhancement stages, respectively.
For the consistency stage, the training data needs to contain realistic motion while maintaining reasonable image quality. To this end, we select 44,162 high-quality frames from the REDS dataset nah2019REDS , which offers professionally captured sequences with rich dynamics, and a curated set of videos wu2025insvie from Pexels [^1] https://www.pexels.com/, chosen based on aesthetic and temporal smoothness criteria. These sequences provide necessary temporal priors for learning degradation-robust representations.
For the enhancement stage, the training data should prioritize visual quality. Thus, we select the LSDIR li2023lsdir dataset, known for its rich textures and fine-grained details. To preserve the learned consistency modeling capability and enable the optical flow regularization among frames, we generate simulated video sequences based on LSDIR. Specifically, for each ground-truth image in LSDIR, we apply random pixel-level translations to it to generate multiple shifted images.
The data in both stages are degraded using the RealESRGAN wang2021realesrgan degradation pipeline. We apply identical degradation parameters across frames within the same video, while using random parameters for different video sequences.
Testing Datasets. We evaluate our method on both synthetic and real-world datasets, including UDM10 yi2019UDM10 , SPMCS tao2017SPMCS , RealVSR yang2021realvsr , and VideoLQ chan2022realbasicvsr . Among them, UDM10 contains 10 sequences, each having 32 frames. SPMCS contains 30 sequences, each having 31 frames. RealVSR contains 50 real-world sequences, each having 50 frames. VideoLQ contains 50 real-world sequences with complex degradations. For the synthetic dataset (UDM10 and SPMCS), we synthesize LQ-HQ pairs following the same degradation pipeline in training. For real-world datasets (RealVSR and VideoLQ), we directly adopt the given LQ-HQ pairs.
Evaluation Metrics. A set of full-reference and no-reference metrics are selected to evaluate different real-world VSR methods. The full-reference metrics include PSNR and SSIM, and perceptual quality with LPIPS zhang2018lpips and DISTS ding2020dists . No-reference quality assessment involves MUSIQ ke2021musiq , MANIQA yang2022maniqa , CLIPIQA wang2023clipiqa , and the video quality assessment metric DOVER wu2023dover . Compared to Real-ISR, Real-VSR places greater emphasis on temporal consistency. Following prior works yang2024motion ; zhang2024tmp , we use the average warping error $E^{*}_{warp}$ to quantitatively assess temporal consistency: $E^{*}_{warp}=\frac{1}{N-1}\sum_{i=1}^{N-1}||I_{i+1}^{HQ}-F_{wp}(I_{i}^{HQ})||_{1}$. For the test datasets with GT, optical flow in $F_{wp}$ is estimated from GT frames. For real-world datasets without GT (e.g., VideoLQ test set), we use the flow estimated from predicted frames.
4.2 Experimental Results
To demonstrate the effectiveness of our DLoRAL algorithm, we compare it with seven represnetative and state-of-the-art methods, including three Real-ISR models (RealESRGAN wang2021realesrgan , StableSR wang2024stablesr , and the one-step model OSEDiff wu2024osediff ), a discriminative VSR model (RealBasicVSR chan2022realbasicvsr ), and three diffusion-based VSR models (Upscale-A-Video zhou2024upscale , MGLD-VSR yang2024motion and STAR xie2025star ).
Datasets Metrics Real-ISR Methods Real-VSR Methods
RealESRGSN StableSR OSEDiff RealBasicVSR Upscale-A-Video MGLD STAR DLoRAL
UDM10 PSNR $\uparrow$ 21.345 22.042 23.761 24.334 22.364 24.192 24.451 23.975
SSIM $\uparrow$ 0.565 0.568 0.696 0.723 0.584 0.685 0.714 0.710
LPIPS $\downarrow$ 0.451 0.455 0.367 0.363 0.410 0.335 0.417 0.327
DISTS $\downarrow$ 0.175 0.185 0.175 0.204 0.198 0.176 0.230 0.179
BRISQUE $\downarrow$ 29.843 26.310 20.718 14.129 17.607 22.701 36.910 16.250
MUSIQ $\uparrow$ 49.838 47.805 63.146 62.360 61.046 61.309 40.789 65.620
CLIPIQA $\uparrow$ 0.474 0.445 0.574 0.474 0.445 0.453 0.267 0.652
MANIQA $\uparrow$ 0.330 0.319 0.334 0.330 0.318 0.291 0.244 0.373
$E^{*}_{\text{warp}}$downarrow 7.580 8.440 5.220 4.670 5.790 4.610 3.510 4.720
DOVER $\uparrow$ 36.860 30.470 48.404 37.572 37.694 40.045 30.384 42.871
SPMCS PSNR $\uparrow$ 21.660 19.260 20.650 21.580 19.030 21.260 20.730 21.240
SSIM $\uparrow$ 0.569 0.585 0.696 0.545 0.386 0.515 0.489 0.524
LPIPS $\downarrow$ 0.444 0.432 0.354 0.404 0.485 0.384 0.606 0.375
DISTS $\downarrow$ 0.246 0.235 0.229 0.237 0.274 0.234 0.342 0.222
BRISQUE $\downarrow$ 25.240 26.310 19.471 12.048 19.784 23.184 27.902 11.030
MUSIQ $\uparrow$ 53.221 47.805 64.619 66.683 66.912 65.079 33.247 67.390
CLIPIQA $\uparrow$ 0.515 0.445 0.526 0.515 0.517 0.437 0.240 0.581
MANIQA $\uparrow$ 0.308 0.319 0.308 0.308 0.443 0.312 0.237 0.340
$E^{*}_{\text{warp}}$downarrow 7.570 8.430 7.500 5.400 7.570 4.410 4.080 6.250
DOVER $\uparrow$ 32.151 30.470 40.160 30.953 32.151 31.118 17.220 34.895
RealVSR PSNR $\uparrow$ 21.340 18.950 19.920 22.270 20.060 21.120 15.080 20.360
SSIM $\uparrow$ 0.565 0.583 0.588 0.720 0.591 0.646 0.433 0.606
LPIPS $\downarrow$ 0.451 0.225 0.282 0.193 0.263 0.219 0.409 0.242
DISTS $\downarrow$ 0.175 0.154 0.164 0.160 0.158 0.151 0.279 0.150
BRISQUE $\downarrow$ 29.843 28.250 31.794 30.362 25.476 39.082 62.750 27.893
MUSIQ $\uparrow$ 49.838 69.962 64.101 71.413 67.714 70.734 67.947 70.908
CLIPIQA $\uparrow$ 0.474 0.612 0.546 0.370 0.436 0.530 0.532 0.617
MANIQA $\uparrow$ 0.330 0.345 0.341 0.384 0.414 0.496 0.438 0.386
$E^{*}_{\text{warp}}$downarrow 17.580 25.010 18.300 18.720 18.200 19.210 24.600 17.300
DOVER $\uparrow$ 36.860 46.846 42.138 46.439 36.136 42.044 30.214 49.646
VideoLQ BRISQUE $\downarrow$ 29.605 22.337 26.403 24.790 25.101 29.606 42.582 23.039
MUSIQ $\uparrow$ 53.138 52.975 58.959 59.475 57.489 53.092 49.305 63.846
CLIPIQA $\uparrow$ 0.334 0.478 0.499 0.393 0.377 0.315 0.333 0.567
MANIQA $\uparrow$ 0.232 0.278 0.254 0.312 0.328 0.254 0.268 0.344
$E^{*}_{\text{warp}}$downarrow 7.580 8.430 8.406 8.108 7.586 7.409 7.280 7.897
DOVER $\uparrow$ 28.400 30.470 37.580 34.772 36.860 31.899 29.400 38.505
Table 1: Comparison of various Real-ISR and Real-VSR methods across different datasets. The best and second best results of each metric are highlighted in red and blue, respectively.
Quantitative Comparison. We show the quantitative comparison on both synthetic and real-world video benchmarks in Tab 1, from which several key observations can be made. First, non-diffusion-based methods (e.g., RealESRGAN and RealBasicVSR) perform worse than diffusion-based methods on no-reference perceptual quality metrics, such as MUSIQ and CLIPIQA, mainly because they lack the strong image priors provided by pre-trained SD models, leading to over-smoothed results. Second, SD-based Real-ISR methods (StableSR and OSEDiff) can achieve comparable or even better perceptual quality scores than existing Real-VSR methods. In particular, OSEDiff achieves the best DOVER scores on both the UDM10 and SPMCS datasets. However, its warping error evaluated by $E^{*}_{warp}$ is worse. This is because the Real-ISR methods generate details for each frame without considering the inter-frame consistency. Finally, compared to the existing Real-VSR methods, our DLoRAL consistently ranks first or second across a range of perceptual quality metrics, including LPIPS, DISTS, MUSIQ, CLIPIQA, MANIQA, and DOVER, demonstrating its strong alignment with human perception. At the same time, DLoRAL does not compromise temporal consistency, as evidenced by its superior performance on the $E^{*}_{warp}$ metric. For example, on the RealVSR dataset, DLoRAL achieves state-of-the-art results in DISTS, CLIPIQA, and DOVER, while ranking among the top in $E^{*}_{warp}$, highlighting its ability to produce visually pleasing and temporally coherent outputs.
It should be mentioned that although $E^{*}_{warp}$ is widely used to assess temporal consistency, it does not correlate well with human perception. For example, blurry Real-VSR outputs can achieve lower warping errors but exhibit poorer visual quality. DLoRAL may report slightly larger $E^{*}_{warp}$ values than some methods (e.g., STAR), but this is because DLoRAL better preserves fine details that the warping error metric tends to penalize.
Qualitative Comparison. To further demonstrate the effectiveness of DLoRAL, we visualize the Real-VSR results in Fig. 3. One can see that DLoRAL can remove complex spatial-variant degradations and generate realistic details, significantly outperforming other Real-VSR models. Specifically, for the severely degraded facial region (first row), RealBasicVSR fails to reconstruct the facial structural, and Upscale-A-Video and STAR lose facial details. MGLD produces sharper outputs, but suffers from severe structural distortions, particularly around the eye regions. In contrast, DLoRAL successfully recovers fine facial features while maintaining structural integrity. The second row highlights the performance in texture reconstruction, where our method restores sharper and more legible texture patterns compared to the blurry or distorted outputs from other algorithms.
To better compare the consistency, we plot the temporal profiles of the VSR results produced by competing methods Fig. 4. Real-ISR approaches such as StableSR and OSEDiff restore sharper details but suffer from severe temporal instability, as shown by the erratic fluctuations in their profiles, resulting in unpleasant flickering that harms the video quality. On the other hand, while existing Real-VSR methods can offer better temporal consistency than Real-ISR methods, this comes at the cost of blurred details (see the results of Upscale-A-Video, MGLD and STAR in the left case of Fig. 4) or intra-frame artifacts (see the results of RealBasicVSR, Upscale-A-Video and STAR in the right case of Fig. 4). In comparison, our DLoRAL produces smooth and stable transitions across frames, as reflected by its consistent temporal profiles. This qualitative evidence aligns with our quantitative results, demonstrating DLoRAL’s ability to preserve fine visual details while ensuring natural temporal consistency. More visual comparisons can be found in the Appendix.
Figure 3: Qualitative comparison of VSR models on real-world VideoLQ dataset.
Figure 4: Temporal profiles of competing Real-ISR and Real-VSR methods.
Real-ISR Methods Real-VSR Methods
StableSR OSEDiff Upscale-A-Video MGLD STAR DLoRAL
Inference Step 200 1 30 50 15 1
Inference Time (s/50 frames) 32800 340 3640 4146 2830 346
# Total Param (M) 1150 1294 14442 1430 2492 1300
Table 2: Complexity comparison among different methods. All methods are evaluated using 50 $512\times 512$ frames for the $\times$4 VSR task. Inference time is measured on an A100 GPU and includes the entire pipeline: data loading, processing, and result storage.
Complexity Comparison. We compare the inference steps, model size, and inference time of competing diffusion-based models in Tab. 2. The inference time of the whole pipeline (including data loading, data processing, and result storage) is reported, which is measured on the $\times$4 VSR task with 50 frames of $512\times 512$ LQ images on a single NVIDIA A100 80G GPU. Compared with Real-ISR methods, DLoRAL (346s) achieves a strong balance between quality and complexity, delivering superior visual quality and temporal consistency while maintaining a similar speed to OSEDiff (340s). Among the Real-VSR methods, DLoRAL achieves the fastest inference time and the lowest parameter count, benefiting from its efficient one-step design. Specifically, DLoRAL is more 10$\times$ faster than Upscale-A-Video, MGLD, and 8$\times$ faster than STAR, while maintaining superior visual quality.
User Study. We also conduct a user study to further examine the effectiveness of DLoRAL in comparison with existing RealVSR methods. We invited ten volunteers to participate in a user study. Our DLoRAL method was compared with the other three diffusion-based Real-VSR methods: Upscale-A-Video zhou2024upscale , MGLD yang2024motion and STAR xie2025star . We randomly selected 12 real-world LQ videos with complex degradations and motions from the VideoLQ dataset chan2022realbasicvsr , whose scenes are shown in Fig. 5(a). Each LQ video and its corresponding HQ videos generated by the competing Real-VSR methods were presented to the participants who were asked to select the best HQ result by considering two equally weighted factors: the perceptual quality and temporal consistency of the video.
The results of the user study are shown in Fig. 5(b). DLoRAL received 93 votes, significantly outperforming the other methods, with MGLD, STAR, and Upscale-A-Video receiving 14, 8, and 5 votes, respectively. This overwhelming preference for DLoRAL highlights its effectiveness in addressing the challenges of real-world video restoration. Note that the selected videos include a variety of motion scenarios. In scenarios with complex motion, DLoRAL is able to achieve superior visual quality while maintaining temporal consistency comparable to other methods. In relatively static scenes, DLoRAL demonstrates stable temporal consistency along with equally sharp and clear visual quality.
Figure 5: LQ videos used in our user study and the voting results.
5 Conclusion
We proposed DLoRAL to achieve temporally consistent and detail-rich Real-VSR results. To effectively extract degradation-robust temporal priors from low-quality input videos while enhancing details without compromising these priors, we first developed a CFR module and a consistency-LoRA to generate robust temporal representations, and then developed a detail-LoRA to enhance spatial details. We optimized these two objectives alternatively and iteratively, where the results of the previous stage served as an anchor to provide priors for the next stage. The resulting DLoRAL model demonstrated significantly superior performance to previous Real-VSR methods, achieving rich spatial details without compromising the temporal coherence.
Limitations. Despite its strong performance, DLoRAL still has certain limitations. First, since it inherits the 8×8$\times$8 × downsampling VAE from SD, DLoRAL faces difficulties in restoring very fine-scale details such as small texts. Second, this heavy compression of VAE may disrupt temporal coherence, making it harder to extract robust consistency priors. A VAE specifically designed for Real-VSR tasks could help to address these issues. We leave this challenge for future investigation.
Appendix A Appendix
In the appendix, we provide the following materials:
  • Training algorithm of DLoRAL (referring to Section 3.2 in the main paper);
  • Ablation studies on training strategies in DLoRAL (referring to Section 3.2 in the main paper);
  • Results of user study (referring to Section 4.2 in the main paper);
  • More real-world visual comparisons under scaling factor 4×4$\times$4 × (referring to Section 4.2 in the main paper);
  • Video demonstration;
  • Broader impacts.
A.1 Training algorithm of DLoRAL
Input: Training datasets $\mathcal{S}_{cons}$ and $\mathcal{S}_{enh}$, pretrained SD with VAE encoder $E_{\phi}$, diffusion UNet $\epsilon_{\phi}$, and VAE decoder $D_{\phi}$, cross-frame retrieval $R$, prompt extractor $Y$, training iteration $N$, consistency step $N_{cons}$, enhancement step $N_{enh}$.
1
2Initialize two LoRA modules parameterized by $\theta_{1}$ and $\theta_{2}$ within diffusion UNet $\epsilon_{\phi}$, denoted as $\epsilon_{\{\phi,\theta_{1},\theta_{2}\}}$
3Initialize cross-frame retrieval module $R$ parameterized by $\mu$, denoted as $R_{\mu}$
4Initialize $N_{cycle}\leftarrow N_{cons}+N_{enh}$
5for $i\leftarrow 1$ to $N$ do
6       if $(i-1)\bmod N_{cycle}<N_{cons}$ then
             /* Consistency Stage */
7             Sample $\boldsymbol{x}_{L},\boldsymbol{x}_{H}$ from $\mathcal{S}_{cons}$
            /* Network forward */
8             $c_{y}\leftarrow\emph{Y}(\boldsymbol{x}_{L})$
9            $\boldsymbol{z}_{L}\leftarrow\emph{E}_{\phi}(\boldsymbol{x}_{L})$
10            $\hat{\boldsymbol{z}}_{H}\leftarrow\boldsymbol{\epsilon}_{\{\phi,\theta_{1},\theta_{2}\}}(R_{\mu}(\boldsymbol{z}_{L});c_{y})$
11            $\hat{\boldsymbol{x}}_{H}\leftarrow\emph{D}_{\phi}(\hat{\boldsymbol{z}}_{H})$
            /* Compute consistency loss group */
12             $\nabla_{\theta_{1}}\mathcal{L}_{\mathrm{data}}\leftarrow\left[\mathcal{L}_{\mathrm{cons}}(\hat{\boldsymbol{x}}_{H},\boldsymbol{x}_{H})\right]\frac{\partial\hat{\boldsymbol{x}}_{H}}{\partial\theta_{1}}$
13            $\nabla_{\mu}\mathcal{L}_{\mathrm{data}}\leftarrow\left[\mathcal{L}_{\mathrm{cons}}(\hat{\boldsymbol{x}}_{H},\boldsymbol{x}_{H})\right]\frac{\partial\hat{\boldsymbol{x}}_{H}}{\partial\mu}$
            /* Update parameters */
14             Update $\theta_{1}$ and $\mu$ with $\mathcal{L}_{\mathrm{data}}$
15       end if
16      else
             /* Enhancement Stage */
17             Sample $\boldsymbol{x}_{L},\boldsymbol{x}_{H}$ from $\mathcal{S}_{enh}$;
            /* Network forward */
18             $c_{y}\leftarrow\emph{Y}(\boldsymbol{x}_{L})$
19            $\boldsymbol{z}_{L}\leftarrow\emph{E}_{\phi}(\boldsymbol{x}_{L})$
20            $\hat{\boldsymbol{z}}_{H}\leftarrow\boldsymbol{\epsilon}_{\{\phi,\theta_{1},\theta_{2}\}}(R_{\mu}(\boldsymbol{z}_{L});c_{y})$
21            $\hat{\boldsymbol{x}}_{H}\leftarrow\emph{D}_{\phi}(\hat{\boldsymbol{z}}_{H})$
            /* Compute enhancement loss group */
22             $\nabla_{\theta_{2}}\mathcal{L}_{\mathrm{data}}\leftarrow\left[\mathcal{L}_{\mathrm{enh}}(\hat{\boldsymbol{x}}_{H},\boldsymbol{x}_{H})\right]\frac{\partial\hat{\boldsymbol{x}}_{H}}{\partial\theta_{2}}$
            /* Update parameters */
23             Update $\theta_{2}$ with $\mathcal{L}_{\mathrm{data}}$
24       end if
25
26 end for
Output: Generator $\emph{G}_{\theta}$ including VAE encoder $E_{\phi}$, cross-frame retrieval $R_{\mu}$, latent diffusion UNet $\boldsymbol{\epsilon}_{\{\phi,\theta_{1},\theta_{2}\}}$ and VAE decoder $D_{\phi}$
Algorithm 1 Training Algorithm of DLoRAL
A.2 Ablation studies on training strategies in DLoRAL
Effectiveness of DLoRAL strategy. To validate the effectiveness of our proposed DLoRAL strategy, we first conduct ablation studies by testing three configurations: (a) using a single parameter space, where optimization is guided by consistency losses and enhancement losses simultaneously; (b) using a single parameter space with iterative optimization guided by consistency losses and enhancement losses; and (c) using a dual parameter space (i.e., two LoRA modules), where optimization is guided by consistency losses and enhancement losses iteratively (referring to the settings in Section 3.3 of the main paper). To evaluate the video quality (measured by MUSIQ ke2021musiq ) and temporal consistency (measured by warping error yang2024motion ) of the models at different training iterations, we select four representative sequences with diverse scenes and motions from the VideoLQ dataset, denoted as VideoLQ4222Specifically, VideoLQ4 contains the 013, 015, 020, and 041 clips, each consisting of 100 frames.. The results are shown in Fig. A1.
For the single parameter space optimized jointly, as illustrated in Fig. A1(a), the final performance is characterized by a relatively low warping error but poor image quality. During training, the decline in warping error is consistently accompanied by a degradation in image quality. Eventually, both warping error and image quality reach a bottleneck, highlighting the inherent difficulty of optimizing both objectives simultaneously. In contrast, the dual-stage training strategy, as illustrated in Fig. A1(c), achieves significantly better image quality while maintaining a comparable level of warping error. Although some conflicts between consistency and image quality remain in each stage, these conflicts diminish with iterations, as evidenced by the reduced warping error fluctuation (from  0.0013 in the first stage to <0.0005 in the second) and smoother MUSIQ improvement (stabilizing around MUSIQ = 66). These results demonstrate that the dual-stage strategy effectively mitigates conflicts and achieves a better balance between the two objectives.
Joint optimization guided by enhancement and consistency losses struggles to optimize both metrics simultaneously. To ensure a fair comparison, we also use a single parameter space with iterative optimization guided by consistency and enhancement losses, as illustrated in Fig. A1(b). The improvement in image quality is obtained at a price of significant increase in warping error, and vice versa. In addition, the similar performance between the first (i.e., 0$\sim$5000 iterations) and second consistency stages (i.e., 10000$\sim$15000) implies that the model may not fully utilize the priors from the intervening enhancement stage to improve the subsequent consistency stage. In contrast, DLoRAL, as shown in Fig. A1(c), leverages the priors learned from the previous stage as anchors, enabling effective information transfer and gradually mitigating conflicts. This allows DLoRAL to achieve significantly higher image quality while maintaining comparable warping error.
Figure A1: Performance (MUSIQ ke2021musiq and Warping Error yang2024motion ) comparisons of different training strategies on the VideoLQ4 test dataset. (a) Joint optimization in a single parameter space struggles to balance image quality and temporal consistency, with both metrics reaching a bottleneck. (b) Iterative optimization fails to transfer priors effectively, leading to limited improvement. (c) In contrast, the dual-stage strategy mitigates conflicts and achieves superior image quality while maintaining comparable warping error. (d) Numeric results of the last iteration.
Enhancement-stage first vs. consistency-stage first. For dual-stage training, an intuitive question arises: Should the first stage focus on detail enhancement or temporal consistency? We conduct experiments to answer this question with two options: an enhancement-first alternating strategy and a consistency-first alternating strategy. The quantitative and qualitative results are presented in Fig. A2 and Tab. A1.
As shown in Tab.A1 and Fig.A2, the enhancement-first model achieves comparable continuity but exhibits less details. This suggests that, although both configurations adopt the dual-stage alternating training strategy, the choice of the first stage impacts the final results. When the enhancement stage is used as the first stage, the model focuses on generating high-resolution details before optimizing the inter-frame warping error. However, the priors learned during the enhancement stage, which focus on enhancing fine details, cannot be directly utilized as priors by the consistency stage, leading to certain degree of misalignment between the two stages. The misalignment consequently leads to poor image quality, as shown in Fig.A2. In contrast, when the consistency stage precedes the enhancement stage, it extracts temporal priors from degraded frames. The temporal priors serve as a foundation for the subsequent enhancement stage, leading to detail-rich and consistent outputs.
[…]
Enhancement-first 46.008 49.484 0.299 0.258 7.040 25.175
Consistency-first 23.039 63.846 0.567 0.344 7.497 38.505
Table A1: Comparison of enhancement-first strategy and consistency-first strategy. The experiments are conducted on the VideoLQ dataset.
Figure A2: Visual comparisons of enhancement-first strategy and consistency-first strategy. The experiments are conducted on the VideoLQ dataset.
A.3 More visual comparisons
Fig. A3 provides more visual comparisons between DLoRAL and other diffusion-based Real-VSR methods. DLoRAL achieves better results in preserving complex textures such as animal fur, plants, and buildings.
Figure A3: More Visual Comparisons of diffusion-based Real-VSR models on real-world and synthetic datasets. Please zoom in for a better view.
A.4 Video demonstration
We provide a demonstration video (DLoRAL.mp4) at https://drive.google.com/file/d/1mwNRKo7SaJkezjptE4PctfKjx4VGNJNc/view?usp=sharing to illustrate the effectiveness of our method on real-world video inputs, particularly highlighting its temporal coherence and high-fidelity detail preservation. Note that the videos are compressed and the original results exhibit higher visual quality. To ensure optimal visual perception and fair comparison of fine structures, please view the videos at 1080p resolution.
A.5 Broader impacts
This paper presents an exploratory study on real-world video super-resolution using one-step diffusion model. The main academic contribution is to advance video super-resolution research and inspire new methods in related fields. In terms of broader societal impact, the proposed method can significantly improve the quality of videos captured by cameras and mobile devices, with potential benefits for areas such as remote education and digital content creation. While higher-quality media may increase data storage demands, these challenges are manageable and outweighed by the advantages of enhanced video resolution.